# SeCon-RAG: A Two-Stage Semantic Filtering and Conflict-Free Framework for Trustworthy RAG

### v1

@@ 4 The Proposed Defense Method for Corpus Poisoning Attacks
To protect Retrieval-Augmented Generation systems from corpus poisoning attacks, we propose SeCon-RAG, a robust two-stage filtering framework designed to detect and suppress poisoning documents. The first stage eliminates poisoned content statistically and semantically, while the second stage ensures factual consistency from a semantic reasoning perspective. This design ensures robustness without unnecessary knowledge loss. To enable fine-grained semantic understanding and aid in the detection of potentially poisoned content, we propose Entity-Intent-Relation Extractor in section 4.1, a semantic structure extraction module that serves as the foundation for our two-stage filtering framework. Before the retrieval stage, we propose Semantic and Cluster-Based Filtering shown in section 4.2 creates a semantic graph from the information extracted by EIRE, allowing for dual-channel filtering based on both clustering structure and semantic relevance. During the inference stage, we introduce the Conflict-Aware Filtering module shown in section 4.3. CAF performs cross-source semantic consistency checks using both EIRE on the retrieved content and the model’s internal knowledge representations. Figure 1 shows an overview of the full Secon-RAG framework. The appendix A shows the pseudocode for the overall algorithm.
@@ 4.1 EIRE: Entity-Intent-Relation Extractor
To extract these components, EIRE employs a prompt-based large language model. Given a document $d$, we create structured prompts that direct the LLM to generate a structured triple. Appendix A.2.5 contains an example of the prompt and its output. For a document $d$, EIRE generates a structured triple $({E}_{d},{I}_{d},{R}_{d})$, where ${E}_{d}$ is the set of extracted entities, ${I}_{d}$ is the identified intent, and ${R}_{d}$ is the set of semantic relations between entities. By grounding document analysis in interpretable semantic frames, EIRE provides a robust and explainable foundation for downstream filtering.
@@ 4.2.1 Clustering-Based Filtering
Adversarially generated poisoning documents often exhibit highly similar phrasing or templated structures, especially when crafted to target the same query. As a result, they naturally form tight clusters in the embedding space zhou2025trustrag . To mitigate this, we first apply a clustering-based filter to detect poisoning document groups. Given a potentially poisoned corpus $\mathcal{D}^{\prime}=\mathcal{D}\cup\Gamma$, we embed each document $d\in\mathcal{D}^{\prime}$ into the vector representation $m(d)$ and apply K-means clustering to obtain $K$ clusters $C=\{c_{1},\dots,c_{K}\}$, each with centroid $\mu_{i}=\frac{1}{|c_{i}|}\sum_{d_{j}\in c_{i}}m(d_{j}),\quad\bigcup_{i=1}^{K}c_{i}=\mathcal{D}^{\prime}$ (na2010research, ). We then define the filtered set as:
@@ 4.2.2 Semantic Graph-Based Filtering by EIRE
To use these structural properties, we first construct a set of semantic graphs $G_{\text{cor}}=\{G_{d_{\text{cor},i}}\}$ from a small collection of verified correct documents $D_{\text{cor}}=\{d_{\text{cor},i}\}$ and use the semantic graphs $G_{\text{cor}}$ as a benchmark. $G_{\text{cor}}$ used for semantic reference are a small set of samples chosen manually from the dataset. For any candidate document $d\in\mathcal{D}^{\prime}$, we generate its semantic graph $G_{d}$ using EIRE and assess its similarity to the $G_{\text{cor}}$. Rather than relying on rigid graph similarity metrics, we employ large language models’ semantic reasoning capabilities to compare graph structures in a more flexible and context-aware way. For any candidate document $d$, we compute its semantic similarity score $ssG$ by comparing generate $G_{d}$ to $G_{d_{cor,i}}$ using a prompt-based LLM as shown in appendix A.2.5:
@@ 4.3 Conflict-Aware Filtering (CAF)
For each candidate document $d\in\mathcal{D}_{k}(q)$, CAF generates structured semantic information using EIRE, which is divided into three components: Entities capable of determining whether facts align with the model’s internal knowledge; Intent to evaluate query alignment; Relations that can evaluate logical coherence across documents and detect contradictions or omissions. In the final inference process, we prompt the LLMs as shown in appendix A.2.5 to determine which information from the retrieve documents is reliable from three dimensions using the semantic information extracted by EIRE:
@@ 5.1 Setup
This section describes the experimental setup. We evaluate the effectiveness and robustness of SeConRAG under various adversarial scenarios. All reported results are averages of multiple runs.
@@
Model. We tested five LLMs from both open and closed source families: Mistral-12B, Qwen-7B, LLaMA-3.1-8B, GPT-4o, and DeepSeek-R1. RAG backbones are maintained in accordance with the corresponding LLMs. Appendix  A.2.5 provides detailed prompts for EIRE, semantic similarity, and CAF modules. All experiments are carried out using NVIDIA A100-SXM4-40GB GPUs.
@@ 5.2 Main Results
We evaluate the robustness and effectiveness of SeconRAG against four representative Retrieval-Augmented Generation (RAG) defense baselines VanillaRAG, InstructRAGwei2024instructrag , AstuteRAGwang2024astute , and TrustRAGzhou2025trustrag , across three datasets and five LLMs. To verify that our method can handle both high and low poisoning rates simultaneously, each model is tested under four settings: Clean, Prompt Injection Attack, PoisonedRAG-20%, and PoisonedRAG-100%, with results reported in terms of Accuracy and Attack Success Rate. Table 1 summarizes the key findings. More detailed PoisonedRAG results at different poisoning levels can be found in Appendix A.2.1.
@@
Model Method HotpotQA Yang2018hotpotqa NQ Kwiatkowski2019natural MS-MARCO bajaj2016ms
@@
Mistral-12B aydin2025generative VanillaRAG 51.0 / 40.0 0.9 / 98.2 38.2 / 58.0 75.0 47.6 / 37.5 8.2 / 90.9 38.2 / 48.2 68.0 54.5 / 43.6 9.1 / 89.1 50.0 / 45.5 84.0
InstructRAG wei2024instructrag 50.0 / 43.6 13.6 / 83.5 45.5 / 49.1 75.0 48.2 / 43.2 13.6 / 82.7 51.8 / 40.0 66.0 64.5 / 33.2 15.5 / 78.2 57.3 / 36.4 81.0
ASTUTERAG wang2024astute 68.2 / 17.3 32.7 / 61.1 65.9 / 21.8 76.0 64.5 / 10.0 43.6 / 38.2 67.7 / 11.8 70.0 75.9 / 17.3 32.7 / 58.2 73.6 / 18.8 81.0
TrustRAG zhou2025trustrag 75.5 / 1.4 75.5 / 3.6 71.8 / 14.5 81.0 68.2 / 0.5 62.7 / 1.8 66.4 / 13.6 73.0 90.9 / 0.0 91.8 / 0.0 87.3 / 11.8 85.0
@@
Qwen-7B hui2024qwen2 VanillaRAG 34.0 / 60.9 1.8 / 98.2 32.7 / 65.5 67.0 28.2 / 67.3 5.5 / 93.6 39.1 /51.8 56.0 36.4 / 60.9 10.0 / 87.3 43.6 / 46.4 75.0
InstructRAG wei2024instructrag 58.2 / 38.2 24.5 / 76.4 45.5 / 51.8 67.0 52.7 / 45.5 25.5 / 76.4 47.3 / 47.3 64.0 61.8 / 36.4 43.6 / 57.8 49.1 / 45.5 75.0
ASTUTERAG wang2024astute 51.8 / 29.1 45.5 / 44.1 58.6 / 25.4 65.0 56.4 / 17.3 42.3 / 53.2 60.5 / 17.3 68.0 44.5 / 45.5 42.3 / 54.5 65.5 / 20.0 74.0
TrustRAG zhou2025trustrag 62.7 / 0.6 58.2 / 2.7 58.2 / 26.4 73.0 67.3 / 0.6 60.0 / 2.7 64.5 / 24.5 67.0 68.2 / 1.8 64.5 / 11.8 66.4 / 22.7 78.0
@@
LLaMA-3.1-8B dubey2024llama VanillaRAG 31.8 / 62.7 4.5 / 96.4 36.4 / 57.3 70.0 38.2 / 54.5 10.9 / 88.2 41.8 / 52.7 70.0 34.5 / 63.6 9.1 / 88.2 54.5 / 40.9 83.0
InstructRAG wei2024instructrag 61.8 / 30.0 27.3 / 71.8 47.3 / 50.0 76.0 67.3 / 24.1 32.7 / 67.3 56.4 / 34.5 70.0 68.2 / 26.4 48.5 / 51.8 72.7 / 27.3 81.0
ASTUTERAG wang2024astute 43.6 / 41.8 46.8 / 47.0 65.5 / 20.9 68.0 57.3 / 26.4 58.2 / 31.8 77.5 / 8.2 81.0 59.1 / 39.5 56.8 / 38.6 82.3 / 13.6 89.0
TrustRAG zhou2025trustrag 72.7 / 0.5 67.3 / 3.0 65.5 / 19.1 72.0 84.5 / 0.2 79.1 / 0.0 79.1 / 10.9 84.0 86.4 / 1.5 84.5 / 6.4 85.4 / 9.1 84.0
@@
GPT-4o achiam2023gpt VanillaRAG 57.3 / 40.0 11.9 / 81.8 45.5 / 30.5 81.0 50.9 / 44.3 27.3 / 68.2 52.7 / 31.8 74.0 70.0 / 27.3 30.0 / 64.1 72.3 / 16.4 84.0
InstructRAG wei2024instructrag 59.1 / 37.3 27.3 / 71.8 61.8 / 33.2 84.0 58.2 / 26.5 43.6 / 51.1 66.4 / 25.5 74.0 77.3 / 16.4 50.5 / 42.7 70.9 / 17.3 83.0
ASTUTERAG wang2024astute 72.7 / 14.5 67.3 / 24.1 77.3 / 11.8 81.0 83.6 / 4.5 75.5 / 14.2 79.1 / 4.1 81.0 90.5 / 0.7 76.4 / 15.5 82.7 / 6.4 86.0
TrustRAG zhou2025trustrag 81.8 / 0.3 80.9 / 2.7 79.1 / 6.4 85.0 82.7 / 0.3 80.0 / 0.1 81.8 / 1.0 86.0 89.1 / 1.3 89.1 / 1.8 84.5 / 6.4 88.0
@@
DeepSeek-R1 guo2025deepseek VanillaRAG 59.1 / 32.7 10.0 / 89.1 51.0 / 46.4 81.0 64.3 / 27.3 17.3 / 84.5 51.0 / 43.6 80.0 71.8 / 25.5 11.8 / 81.8 60.5 / 29.1 82.0
InstructRAG wei2024instructrag 61.8 / 34.5 27.3 / 72.7 61.8 / 38.2 80.0 59.1 / 28.2 39.1 / 62.7 65.5 / 32.7 82.0 75.5 / 18.2 51.8 / 47.5 72.7 / 26.4 87.0
ASTUTERAG wang2024astute 73.6 / 14.5 64.5 / 25.5 77.3 / 14.5 79.0 90.0 / 1.8 81.8 / 10.9 89.1 / 0.0 87.0 87.3 / 4.5 85.5 / 8.2 89.1 / 5.5 88.0
TrustRAG zhou2025trustrag 81.8 / 4.5 79.1 / 2.7 85.5 / 10.0 89.0 90.0 / 1.8 88.2 / 0.0 90.0 / 3.6 91.0 93.6 / 1.8 89.1 / 3.6 89.1 / 5.5 91.0
@@ 5.3 Ablation Study
We conduct an ablation study on the Mistral-12B model to evaluate the contributions of SeCon-RAG’s components. We focus on three aspects: (i) the core SCF and CAF modules, (ii) impact of SCF Subcomponents, and (iii) the effect of EIRE and verified Correct Document Set.
@@
5.3.2 Impact of SCF Subcomponents.
To examine whether clustering and semantic filtering are complementary, we test each individually. As shown in Table 6, their combination yields the strongest robustness, achieving lower ASR in several cases, which confirms the necessity of combination.
Table 2: Ablation of SCF components on Mistral-12B .
PIA 100% 20% PIA 100% 20% PIA 100% 20%
5.3.3 Impact of EIRE and the Verified Correct Document Set.
Finally, we assess the standalone effectiveness of the Entity-Intent-Relation Extractor and the verified correct document set ($d_{cor}$). Table  7 in appendix summarizes the results. EIRE improves the fine-grained reasoning capability of both SCF and CAF. With EIRE enabled, the model consistently achieves higher factual accuracy while significantly lowering ASR, especially under high poisoning conditions. Similarly, a small, high-quality $d_{cor}$ set can significantly improve semantic filtering performance and reduce noise from poisoned documents, as well as improve robustness under high-poisoning conditions (e.g., ASR $\rightarrow 0$ on MS-MARCO 100% poisoning).

### v2

@@ 4 The Proposed Defense Method for Corpus Poisoning Attacks
To protect Retrieval-Augmented Generation systems from corpus poisoning attacks, we propose SeCon-RAG, a robust two-stage filtering framework designed to detect and suppress poisoning documents. The first stage eliminates poisoned content statistically and semantically, while the second stage ensures factual consistency from a semantic reasoning perspective. This design ensures robustness without unnecessary knowledge loss. To enable fine-grained semantic understanding and aid in the detection of potentially poisoned content, we propose Entity-Intent-Relation Extractor in section 4.1, a semantic structure extraction module that serves as the foundation for our two-stage filtering framework. Before the retrieval stage, we propose Semantic and Cluster-Based Filtering shown in section 4.2 creates a semantic graph from the information extracted by EIRE, allowing for dual-channel filtering based on both clustering structure and semantic relevance. During the inference stage, we introduce the Conflict-Aware Filtering module shown in section 4.3. CAF performs cross-source semantic consistency checks using both EIRE on the retrieved content and the model’s internal knowledge representations. Figure 1 shows an overview of the full Secon-RAG framework. The appendix A.3 shows the pseudocode for the overall algorithm.
@@ 4.1 EIRE: Entity-Intent-Relation Extractor
To extract these components, EIRE employs a prompt-based large language model. Given a document $d$, we create structured prompts that direct the LLM to generate a structured triple. Appendix A.1 contains an example of the prompt and its output. For a document $d$, EIRE generates a structured triple $({E}_{d},{I}_{d},{R}_{d})$, where ${E}_{d}$ is the set of extracted entities, ${I}_{d}$ is the identified intent, and ${R}_{d}$ is the set of semantic relations between entities. By grounding document analysis in interpretable semantic frames, EIRE provides a robust and explainable foundation for downstream filtering.
@@ 4.2.1 Clustering-Based Filtering
Adversarially generated poisoning documents often exhibit highly similar phrasing or templated structures, especially when crafted to target the same query. As a result, they naturally form tight clusters in the embedding space (zhou2025trustrag, ). To mitigate this, we first apply a clustering-based filter to detect poisoning document groups. Given a potentially poisoned corpus $\mathcal{D}^{\prime}=\mathcal{D}\cup\Gamma$, we embed each document $d\in\mathcal{D}^{\prime}$ into the vector representation $m(d)$ and apply K-means clustering to obtain $K$ clusters $C=\{c_{1},\dots,c_{K}\}$, each with centroid $\mu_{i}=\frac{1}{|c_{i}|}\sum_{d_{j}\in c_{i}}m(d_{j}),\quad\bigcup_{i=1}^{K}c_{i}=\mathcal{D}^{\prime}$ (na2010research, ). We then define the filtered set as:
@@ 4.2.2 Semantic Graph-Based Filtering by EIRE
To use these structural properties, we first construct a set of semantic graphs $G_{\text{cor}}=\{G_{d_{\text{cor},i}}\}$ from a small collection of verified correct documents $D_{\text{cor}}=\{d_{\text{cor},i}\}$ and use the semantic graphs $G_{\text{cor}}$ as a benchmark. $G_{\text{cor}}$ used for semantic reference are a small set of samples chosen manually from the dataset. For any candidate document $d\in\mathcal{D}^{\prime}$, we generate its semantic graph $G_{d}$ using EIRE and assess its similarity to the $G_{\text{cor}}$. Rather than relying on rigid graph similarity metrics, we employ large language models’ semantic reasoning capabilities to compare graph structures in a more flexible and context-aware way. For any candidate document $d$, we compute its semantic similarity score $ssG$ by comparing generate $G_{d}$ to $G_{d_{cor,i}}$ using a prompt-based LLM as shown in appendix A.1:
@@ 4.3 Conflict-Aware Filtering (CAF)
For each candidate document $d\in\mathcal{D}_{k}(q)$, CAF generates structured semantic information using EIRE, which is divided into three components: Entities capable of determining whether facts align with the model’s internal knowledge; Intent to evaluate query alignment; Relations that can evaluate logical coherence across documents and detect contradictions or omissions. In the final inference process, we prompt the LLMs as shown in appendix A.1 to determine which information from the retrieve documents is reliable from three dimensions using the semantic information extracted by EIRE:
@@ 5.1 Setup
This section describes the experimental setup. We evaluate the effectiveness and robustness of SeConRAG under various adversarial scenarios. All reported results are averages of multiple runs with an error of ± 1%.
@@
Model. We tested five LLMs from both open and closed source families: Mistral-12B, Qwen-7B, LLaMA-3.1-8B, GPT-4o, and DeepSeek-R1. RAG backbones are maintained in accordance with the corresponding LLMs. Appendix  A.1 provides detailed prompts for EIRE, semantic similarity, and CAF modules. All experiments are carried out using NVIDIA A100-SXM4-40GB GPUs.
@@ 5.2 Main Results
We evaluate the robustness and effectiveness of SeconRAG against four representative Retrieval-Augmented Generation (RAG) defense baselines VanillaRAG, InstructRAG (wei2024instructrag, ), AstuteRAG (wang2024astute, ), and TrustRAG (zhou2025trustrag, ), across three datasets and five LLMs. To verify that our method can handle both high and low poisoning rates simultaneously, each model is tested under four settings: Clean, Prompt Injection Attack, PoisonedRAG-20%, and PoisonedRAG-100%, with results reported in terms of Accuracy and Attack Success Rate. Table 1 summarizes the key findings. More detailed PoisonedRAG results at different poisoning levels can be found in Appendix A.4.1.
@@
Model Method HotpotQA (Yang2018hotpotqa, ) NQ (Kwiatkowski2019natural, ) MS-MARCO (bajaj2016ms, )
@@
Mistral-12B (aydin2025generative, ) VanillaRAG 51.0 / 40.0 0.9 / 98.2 38.2 / 58.0 75.0 47.6 / 37.5 8.2 / 90.9 38.2 / 48.2 68.0 54.5 / 43.6 9.1 / 89.1 50.0 / 45.5 84.0
InstructRAG (wei2024instructrag, ) 50.0 / 43.6 13.6 / 83.5 45.5 / 49.1 75.0 48.2 / 43.2 13.6 / 82.7 51.8 / 40.0 66.0 64.5 / 33.2 15.5 / 78.2 57.3 / 36.4 81.0
ASTUTERAG (wang2024astute, ) 68.2 / 17.3 32.7 / 61.1 65.9 / 21.8 76.0 64.5 / 10.0 43.6 / 38.2 67.7 / 11.8 70.0 75.9 / 17.3 32.7 / 58.2 73.6 / 18.8 81.0
TrustRAG (zhou2025trustrag, ) 75.5 / 1.4 75.5 / 3.6 71.8 / 14.5 81.0 68.2 / 0.5 62.7 / 1.8 66.4 / 13.6 73.0 90.9 / 0.0 91.8 / 0.0 87.3 / 11.8 85.0
@@
Qwen-7B (hui2024qwen2, ) VanillaRAG 34.0 / 60.9 1.8 / 98.2 32.7 / 65.5 67.0 28.2 / 67.3 5.5 / 93.6 39.1 /51.8 56.0 36.4 / 60.9 10.0 / 87.3 43.6 / 46.4 75.0
InstructRAG (wei2024instructrag, ) 58.2 / 38.2 24.5 / 76.4 45.5 / 51.8 67.0 52.7 / 45.5 25.5 / 76.4 47.3 / 47.3 64.0 61.8 / 36.4 43.6 / 57.8 49.1 / 45.5 75.0
ASTUTERAG (wang2024astute, ) 51.8 / 29.1 45.5 / 44.1 58.6 / 25.4 65.0 56.4 / 17.3 42.3 / 53.2 60.5 / 17.3 68.0 44.5 / 45.5 42.3 / 54.5 65.5 / 20.0 74.0
TrustRAG (zhou2025trustrag, ) 62.7 / 0.6 58.2 / 2.7 58.2 / 26.4 73.0 67.3 / 0.6 60.0 / 2.7 64.5 / 24.5 67.0 68.2 / 1.8 64.5 / 11.8 66.4 / 22.7 78.0
@@
LLaMA-3.1-8B (dubey2024llama, ) VanillaRAG 31.8 / 62.7 4.5 / 96.4 36.4 / 57.3 70.0 38.2 / 54.5 10.9 / 88.2 41.8 / 52.7 70.0 34.5 / 63.6 9.1 / 88.2 54.5 / 40.9 83.0
InstructRAG (wei2024instructrag, ) 61.8 / 30.0 27.3 / 71.8 47.3 / 50.0 76.0 67.3 / 24.1 32.7 / 67.3 56.4 / 34.5 70.0 68.2 / 26.4 48.5 / 51.8 72.7 / 27.3 81.0
ASTUTERAG (wang2024astute, ) 43.6 / 41.8 46.8 / 47.0 65.5 / 20.9 68.0 57.3 / 26.4 58.2 / 31.8 77.5 / 8.2 81.0 59.1 / 39.5 56.8 / 38.6 82.3 / 13.6 89.0
TrustRAG (zhou2025trustrag, ) 72.7 / 0.5 67.3 / 3.0 65.5 / 19.1 72.0 84.5 / 0.2 79.1 / 0.0 79.1 / 10.9 84.0 86.4 / 1.5 84.5 / 6.4 85.4 / 9.1 84.0
@@
GPT-4o (achiam2023gpt, ) VanillaRAG 57.3 / 40.0 11.9 / 81.8 45.5 / 30.5 81.0 50.9 / 44.3 27.3 / 68.2 52.7 / 31.8 74.0 70.0 / 27.3 30.0 / 64.1 72.3 / 16.4 84.0
InstructRAG (wei2024instructrag, ) 59.1 / 37.3 27.3 / 71.8 61.8 / 33.2 84.0 58.2 / 26.5 43.6 / 51.1 66.4 / 25.5 74.0 77.3 / 16.4 50.5 / 42.7 70.9 / 17.3 83.0
ASTUTERAG (wang2024astute, ) 72.7 / 14.5 67.3 / 24.1 77.3 / 11.8 81.0 83.6 / 4.5 75.5 / 14.2 79.1 / 4.1 81.0 90.5 / 0.7 76.4 / 15.5 82.7 / 6.4 86.0
TrustRAG (zhou2025trustrag, ) 81.8 / 0.3 80.9 / 2.7 79.1 / 6.4 85.0 82.7 / 0.3 80.0 / 0.1 81.8 / 1.0 86.0 89.1 / 1.3 89.1 / 1.8 84.5 / 6.4 88.0
@@
DeepSeek-R1 (guo2025deepseek, ) VanillaRAG 59.1 / 32.7 10.0 / 89.1 51.0 / 46.4 81.0 64.3 / 27.3 17.3 / 84.5 51.0 / 43.6 80.0 71.8 / 25.5 11.8 / 81.8 60.5 / 29.1 82.0
InstructRAG (wei2024instructrag, ) 61.8 / 34.5 27.3 / 72.7 61.8 / 38.2 80.0 59.1 / 28.2 39.1 / 62.7 65.5 / 32.7 82.0 75.5 / 18.2 51.8 / 47.5 72.7 / 26.4 87.0
ASTUTERAG (wang2024astute, ) 73.6 / 14.5 64.5 / 25.5 77.3 / 14.5 79.0 90.0 / 1.8 81.8 / 10.9 89.1 / 0.0 87.0 87.3 / 4.5 85.5 / 8.2 89.1 / 5.5 88.0
TrustRAG (zhou2025trustrag, ) 81.8 / 4.5 79.1 / 2.7 85.5 / 10.0 89.0 90.0 / 1.8 88.2 / 0.0 90.0 / 3.6 91.0 93.6 / 1.8 89.1 / 3.6 89.1 / 5.5 91.0
@@ 5.3 Ablation Study
We conduct an ablation study on the Mistral-12B model to evaluate the contributions of SeCon-RAG’s components.We concentrate on two key areas: (i) the core SCF and CAF modules, and (ii) an ablation study that includes the SCF subcomponents, EIRE, and the verified Correct Document Set.
@@
5.3.2 Evaluation of SCF Subcomponents, EIRE, and the Verified Correct Document Set.
To examine whether clustering and semantic filtering are complementary, we test each individually. As shown in Table  5, their combination yields the strongest robustness, achieving lower ASR in several cases, which confirms the necessity of combination.
In addition, we assess the standalone effectiveness of the Entity-Intent-Relation Extractor and the verified correct document set ($d_{cor}$). Table  6 in appendix summarizes the results. EIRE improves the fine-grained reasoning capability of both SCF and CAF. With EIRE enabled, the model consistently achieves higher factual accuracy while significantly lowering ASR, especially under high poisoning conditions. Similarly, a small, high-quality $d_{cor}$ set can significantly improve semantic filtering performance and reduce noise from poisoned documents, as well as improve robustness under high-poisoning conditions (e.g., ASR $\rightarrow 0$ on MS-MARCO 100% poisoning).